# Theia: Distilling Diverse Vision Foundation Models for Robot Learning

### v1

@@ Abstract
Vision-based robot policy learning, which maps visual inputs to actions, necessitates a holistic understanding of diverse visual tasks beyond single-task needs like classification or segmentation. Inspired by this, we introduce Theia, a vision foundation model for robot learning that distills multiple off-the-shelf vision foundation models trained on varied vision tasks. Theia’s rich visual representations encode diverse visual knowledge, enhancing downstream robot learning. Extensive experiments demonstrate that Theia outperforms its teacher models and prior robot learning models using less training data and smaller model sizes. Additionally, we quantify the quality of pre-trained visual representations and hypothesize that higher entropy in feature norm distributions leads to improved robot learning performance. Code and models are available here.
@@ 1 Introduction
Vision-based robot policy learning, which learns action policies from visual inputs, requires strong and diverse visual comprehension. These policies involve many implicit vision tasks such as object recognition and semantic grounding, where off-the-shelf VFMs corresponding to some well-defined tasks can be easily found, but there is no single model for all of vision tasks. Studies have shown that off-the-shelf VFMs such as CLIP [2] usually under-perform relative to visual representation models tailored for specific tasks in robot learning [17, 18, 19, 20, 21]. This fact reveals a gap between the needs of robot learning and the limited visual understanding capabilities of any individual VFM. Prior work on learning foundational visual representation models for robotics has focused primarily on improving training data [17, 21, 20] and designing objective functions [17, 18], rather than enhancing the ability to solve multiple implicit visual understanding tasks.
@@
We introduce Theia, a robot vision foundation model that simultaneously distills off-the-shelf VFMs such as CLIP [2], DINOv2 [7], and ViT [5]. Theia generates rich representations for downstream robot learning, maintaining a comprehensive understanding of many visual sub-problems at the spatial level. Compared to off-the-shelf VFMs [2, 5, 7] and prior works [19, 20], Theia offers both better pre-trained visual representations for higher downstream robot learning performance and reduced computational costs. Furthermore, training Theia only requires ImageNet [23] and about 150 GPU hours, in contrast to prior works which necessitate substantially more compute [19, 20, 17, 18]. To understand what makes a good visual representation for robot learning, we observe multiple factors that relate to the performance of downstream robot learning tasks. We hypothesize that higher entropy in representation norms [24] correlates with improved robot learning performance.
@@ 2.1 Visual Representations for Robot Learning
Visual representations are important for vision-based robot policies to parse high-dimensional visual signals. Visual representation learning can happen at different stages, including pre-training [17, 18, 20, 21], joint-learning with robot tasks [25, 26, 27, 28], or a combination of both using either trainable or frozen visual representations [29, 30]. Off-the-shelf visual encoders [4, 5, 2] can also provide visual representations for robot learning. Additionally, an important factor in training visual representations is the choice of data. ImageNet [23], as suggested by Dasari et al. [21], is a particularly effective pre-training dataset, while video datasets [31, 32] are also widely used. Training objectives and auxiliary tasks for visual representation learning vary, including data augmentation [25, 26], prediction tasks [28], contrastive learning [27, 33, 34], and self-supervised learning [19, 20]. Specifically, to handle invariance and equivariance in visual observations, inductive biases [35, 36] and constraints [37, 38] can be introduced into neural networks to improve visual representation quality. Unlike prior works, we build a robot vision foundation model from the perspective of merging the visual understanding abilities of VFMs via distillation.
@@ 2.2 Vision Foundation Models
Vision foundation models trained on large-scale data exhibit strong task-specific performance and transferability to unseen domains and new tasks. VFMs can focus on single tasks, multiple tasks, or remain task-agnostic at the pre-training stage. For example, ViT [5], DeiT [39], and ConvNeXt [40] are designed for image classification, SAM [10] for semantic segmentation, and Depth-Anything [1] for depth prediction. When combined with Large Language Models (LLMs) [41, 3, 42, 43], these models can solve multiple visual tasks including referring segmentation, visual question answering (VQA), and image editing. Task-agnostic models like the DINO series [6, 7] are trained through self-distillation, while CLIP [2] is trained by aligning image-text pairs. Given the strong generalization capability of VFMs, robot learning should also benefit from leveraging the latent representations of pre-trained VFMs, which is the primary motivation for our research.
@@ 3 Method
Given an input image $\mathbf{x}$, the visual encoder $f(\cdot)$ produces a rich representation $\mathbf{z}=f(\mathbf{x})$ (called the Theia-representation), which is utilized for downstream robot learning tasks. We focus on backbone models that are smaller than typical VFMs, specifically using ViT-Tiny, Small, and Base [5, 39] [^1] For simplicity, we denote DeiT-Tiny and DeiT-Small [39] by ViT-Tiny and ViT-Small. DeiT was introduced by Touvron et al. [39] and the original ViT [5] does not have Tiny and Small models. motivated by the limited computing resources on robotic systems. We use Theia-Tiny (Theia-T), Theia-Small (Theia-S) and Theia-Base (Theia-B) to refer to Theia models using a ViT backbone with the corresponding size. To train the Theia-representation, we perform feature distillation with the help of feature translators, which will be described below.
@@ 3.2 Training
To properly accommodate the different scales of teacher representations, we first perform a normalization step. The teacher representations are normalized over each latent dimension, where mean and variance are calculated from all ImageNet training samples.
@@ 4 Experiments
Figure 3: Simulation and real-world (labeled in blue) tasks used in this work. For simulated environments we show one image per task. For real-world tasks, we show images of key steps throughout the task labeled by numbers. A third-person view image is used to show the Spot setup in Drawer Opening.
@@ 4.2 Simulation Results
We comprehensively evaluate Theia and baseline pre-trained models on the MuJoCo subset of CortexBench [20] to provide an overall assessment of pre-trained visual representations. We consider prior works on visual representations for robot learning, including R3M [17], VIP [18], MVP [19], and VC-1 [20], as well as agglomerative models for vision tasks RADIO and E-RADIO [50], and off-the-shelf vision foundation models ViT [5], DINOv2 [7], and CLIP [2]. All pre-trained representations are frozen in this experiment. Throughout this section, we answer the following questions:
@@
We evaluate Transformer-based models (all models in Figure 4 except R3M and VIP) using either their [CLS] token or spatial tokens for downstream robot learning. To accommodate spatial tokens in the robot policy, we introduce extra shallow CNN layers at the input of the policy network, known as “compression layer” [68, 20]. Figure 4 shows the results of all models we evaluated, clearly showing that for Transformer-based models, providing spatial tokens is consistently better than using the [CLS] token for robot learning. This finding applies to both off-the-shelf VFMs and prior pre-trained representations, including MVP and VC-1.
@@ 4.3 Real World Robot Learning
Based on simulation performance, we test Theia-B and the best-performing baseline models: MVP-L [19], R3M [17], VC-1-L [20], DINOv2-L [7], ViT-H [5], and E-RADIO-L [50] for evaluation on real-world tasks. We employ four tasks (Figure 3): Door Opening, Pick-and-Place, and Toy-Microwave Cooking with a WidowX 250s arm, and Drawer Opening with a Boston Dynamics Spot. We train behavioral cloning policies on top of visual representations using conventional policy networks composed of CNNs and MLPs in the WidowX setup and diffusion policy [69] in the Spot setup. During testing, we vary the robot position for Door Opening and Drawer Opening, randomize the object position for Pick-and-Place, and randomize both the object positions and object types in Toy-Microwave Cooking. Full experimental settings are available in the Appendix.
@@ 5 What Makes Visual Representations Good for Robot Learning?
Darcet et al. [24] analyzed the norm of spatial tokens in vision Transformers and found that high-norm outlier tokens are detrimental to vision task performance. Following this, we investigate whether a similar phenomenon arises in visual representations for robot learning. We inspect the feature norms of Theia with different teacher combinations and baseline models evaluated in Section 4.2, and their corresponding performance on the MuJoCo subset tasks. We sample 1% of the visual observations from the MuJoCo task training set and calculate the L2-norm of each spatial token after encoding. We measure the entropy of the feature norm distribution across all samples and patches per model and use it as a quantitative metric.
We confirm that similar outlier tokens also appear in VC-1 corresponding to the image patches that are not task-relevant, shown in the visualizations of feature norms on the right of Figure 7. In contrast, Theia has very few or no outlier tokens, and the tokens with higher norms are more task-relevant even though Theia-representations are not trained on these robot images. In our quantitative analysis (Figure 7, left), we divide the models into distilled and regular based on the observation that distilled models generally have higher entropy (fewer outliers, Figure 4(c) in Darcet et al. [24]). We find that there is a strong correlation (R=0.943) between entropy and robot learning performance among regular models, and a high correlation (R=0.638) among distilled models. We hypothesize that spatial token representations with high entropy (better feature diversity) encode more information that aids policy learning, while less diverse representations (low entropy) may hinder it. In the Appendix, we discuss the results of other quantitative measurements, including feature similarity and PCA-explained variance ratios, where no strong correlations are found.
@@ D.3.1 WidowX Arm Experiments
We train a behavior-cloning policy for each of the four evaluated setups and for each evaluated baseline (see Table 2 and Figure 3); the training hyperparameters are shown in Table 9.
@@
In this task the robot has to open a fridge door in a toy-kitchen setup; we identify two stages to evaluate the task’s success: Open and Fully Open (see Figure 3). We place the robot in front of the fridge and collect 63 demonstrations to train the behavioral cloning policy. We vary the height (z-axis) of the robot base between 40-46cm, and the position (x-axis, parallel to the toy kitchen) of the robot base. At inference time, we vary the height of the robot base among $\{40,42,43,44,46\}$ cm, and select between 5 randomly-picked positions along the x-axis (for all policies). For each robot position, we evaluate the policy twice, for a total of 50 runs.
@@
In this task, the robot has to pick up a pink cup from a toy-sink and drop it into a drying rack located on the left of the sink. We collected 48 demonstrations to train the policy. During evaluation, we vary the cup’s starting position amongst a total of 10 positions, of which 8 positions are equally distributed about the perimeter, and 2 are in the center of the sink. We also roughly vary the direction of the cup handle towards the left or the right. In total, we evaluate this task for 20 runs. There are two key stages for which we measure the success rate: picking up the cup and successfully releasing it into the drying rack.
@@
In this task, the robot has to pick up an object from within the pot on the stovetop, putting the object into a toy-microwave, and closing the microwave. In each test, we initialize the environment with the microwave door open. In this task, we collected 100 demonstrations across 10 different toy-food objects (10 demonstrations per object) with randomized object positions. During evaluation, we test 40 runs on 10 seen objects (4 runs per object), and 10 runs on 5 unseen objects (2 each), for a total of 50 runs. Furthermore, we vary the position of the pot that holds the object. The longer horizon and the variety of objects in this task make it particularly challenging, so using frozen visual encoders was not effective (0% success rate). However, with fine-tuning, the policies performed reasonably well.
@@ D.3.2 Spot Experiments
In this task, the robot has to open the top drawer of a cabinet. The policy receives color images from a single forward facing camera mounted on the body of the Spot. A fiducial marker is used to enable the robot to walk to a random position and orientation for each trial. The starting locations vary by ±$\pm 5$ cm in x and y and ±$\pm 0.2$ radians in orientation. 50 successful demonstrations were collected using a scripted policy and we evaluate each policy for 20 trials. A trial is considered successful if the drawer is opened at least 10 cm. After each trial, a scripted policy uses the fiducial marker to reset the environment and the robot moves to a new random location.

### v2

@@ Abstract
Vision-based robot policy learning, which maps visual inputs to actions, necessitates a holistic understanding of diverse visual tasks beyond single-task needs like classification or segmentation. Inspired by this, we introduce Theia, a vision foundation model for robot learning that distills multiple off-the-shelf vision foundation models trained on varied vision tasks. Theia’s rich visual representations encode diverse visual knowledge, enhancing downstream robot learning. Extensive experiments demonstrate that Theia outperforms its teacher models and prior robot learning models using less training data and smaller model sizes. Additionally, we quantify the quality of pre-trained visual representations and hypothesize that higher entropy in feature norm distributions leads to improved robot learning performance. Code, models, and demo are available here.
@@ 1 Introduction
Vision-based robot policy learning, which learns action policies from visual inputs, requires strong and diverse visual comprehension. These policies involve many implicit vision tasks such as object recognition and semantic grounding, where off-the-shelf VFMs corresponding to some well-defined tasks can be easily found, but there is no single model for all of vision tasks. Studies have shown that off-the-shelf VFMs such as CLIP [2] usually under-perform relative to visual representation models tailored for specific tasks in robot learning [17, 18, 19, 20, 21]. This fact reveals a gap between the needs of robot learning and the limited visual understanding capabilities of any individual VFM. Our motivation is different than all prior works on learning visual representation models for robotics, where those works focused primarily on improving training data  [17, 21, 20], designing objective functions [17, 18], and directly taking advances of vision architectures [19]. We uniquely focus on improving the visual representation from the angle of solving multiple implicit visual understanding tasks, which will benefit the downstream robot learning.
@@
We introduce Theia, a robot vision foundation model that simultaneously distills off-the-shelf VFMs such as CLIP [2], DINOv2 [7], and ViT [5]. Compared to off-the-shelf VFMs [2, 5, 7] and prior works [19, 20], Theia offers both better pre-trained visual representations for higher downstream robot learning performance and reduced computational costs. Furthermore, training Theia only requires ImageNet [23] and about 150 GPU hours, in contrast to prior works which necessitate substantially more compute [19, 20, 17, 18]. To understand what makes a good visual representation for robot learning, we observe multiple factors that relate to the performance of downstream robot learning tasks. We hypothesize that higher entropy in representation norms [24] correlates with improved robot learning performance.
@@ 2.1 Visual Representations for Robot Learning
Visual representations are important for vision-based robot policies to parse high-dimensional visual signals. Visual representation learning can happen at different stages, including pre-training [17, 18, 20, 21], joint-learning with robot tasks [25, 26, 27, 28], or a combination of both using either trainable or frozen visual representations [29, 30]. Off-the-shelf visual encoders [4, 5, 2, 31, 32] can also provide visual representations for robot learning. Additionally, an important factor in training visual representations is the choice of data. ImageNet [23], as suggested by Dasari et al. [21], is a particularly effective pre-training dataset, while video datasets [33, 34] are also widely used. Training objectives and auxiliary tasks for visual representation learning vary, including data augmentation [25, 26], prediction tasks [28], contrastive learning [27, 35, 36], and self-supervised learning [19, 20]. Specifically, to handle invariance and equivariance in visual observations, inductive biases [37, 38] and constraints [39, 40] can be introduced into neural networks to improve visual representation quality. Unlike prior works, we build a robot vision foundation model from the perspective of merging the visual understanding abilities of VFMs via distillation. Concurrent work OpenVLA [32] uses pre-trained SigLIP [41] and DINOv2 [7] encoders, which suggests the benefit of diverse visual understanding in robot tasks.
@@ 2.2 Vision Foundation Models
Vision foundation models trained on large-scale data exhibit strong task-specific performance and transferability to unseen domains and new tasks. VFMs can focus on single tasks, multiple tasks, or remain task-agnostic at the pre-training stage. For example, ViT [5], DeiT [42], and ConvNeXt [43] are designed for image classification, SAM [10] for semantic segmentation, and Depth-Anything [1] for depth prediction. When combined with Large Language Models (LLMs) [44, 3, 45, 46], these models can solve multiple visual tasks including referring segmentation, visual question answering (VQA), and image editing. Task-agnostic models like the DINO series [6, 7] are trained through self-distillation, while CLIP [2] is trained by aligning image-text pairs. Given the strong generalization capability of VFMs, robot learning should also benefit from leveraging the latent representations of pre-trained VFMs. which is the primary motivation for our research.
@@ 3 Method
Given an input image $\mathbf{x}$, the visual encoder $f(\cdot)$ produces a rich representation $\mathbf{z}=f(\mathbf{x})$ (called the Theia-representation), which is utilized for downstream robot learning tasks. We focus on backbone models that are smaller than typical VFMs, specifically using ViT-Tiny, Small, and Base [5, 42] [^1] For simplicity, we denote DeiT-Tiny and DeiT-Small [42] by ViT-Tiny and ViT-Small. DeiT was introduced by Touvron et al. [42] and the original ViT [5] does not have Tiny and Small models. motivated by the limited computing resources on robotic systems. We use Theia-Tiny (Theia-T), Theia-Small (Theia-S) and Theia-Base (Theia-B) to refer to Theia models using a vision transformer backbone with the corresponding size. To train the Theia-representation, we perform feature distillation with the help of feature translators, which will be described below.
@@ 3.2 Training
To properly accommodate the different scales of teacher representations, we first perform a normalization step. This helps us scale the loss of different teacher features evenly and avoid biasing (collapsing) to a teacher model with extremely larger norms. We perform the normalization on the teacher representations over each latent dimension, where mean and variance are calculated from all ImageNet training samples. Details are in Appendix A.
@@ 4 Experiments
Figure 3: Simulation and real-world (labeled in blue) tasks used in this work. For simulated environments we show one image per task. For real-world tasks, we show images of key steps throughout the task labeled by numbers. A third-person view image shows the setup in Drawer Opening.
@@ 4.2 Simulation Results
We comprehensively evaluate Theia and baseline pre-trained models on the MuJoCo subset of CortexBench [20] to provide an overall assessment of pre-trained visual representations. We consider prior works on visual representations for robot learning, including R3M [17], VIP [18], MVP [19], and VC-1 [20], as well as agglomerative models for vision tasks RADIO and E-RADIO [53], and off-the-shelf vision foundation models ViT [5], DINOv2 [7], and CLIP [2]. We also test a naive concatination of three VFMs (CLIP, ViT, and DINOv2) which are used to train Theia, referred as CDV. All pre-trained representations are frozen in this experiment. Throughout this section, we answer the following questions:
@@
We evaluate Transformer-based models (all models in Figure 4 except R3M and VIP) using either their [CLS] token or spatial tokens [5] for downstream robot learning. To accommodate spatial tokens in the robot policy, we introduce extra shallow CNN layers at the input of the policy network, known as “compression layer” [71, 20]. Figure 4 shows the results of all models we evaluated, clearly showing that for Transformer-based models, providing spatial tokens is consistently better than using the [CLS] token for robot learning. This finding applies to both off-the-shelf VFMs and prior pre-trained representations, including MVP and VC-1.
@@ 4.3 Real World Robot Learning
Based on simulation performance, we test Theia-B and the best-performing baseline models: MVP-L [19], R3M [17], VC-1-L [20], DINOv2-L [7], ViT-H [5], and E-RADIO-L [53] for evaluation on real-world tasks. We employ four tasks (Figure 3): Door Opening, Pick-and-Place, and Toy-Microwave Cooking with a WidowX 250s arm, and Drawer Opening with a Boston Dynamics Spot. We train behavioral cloning policies on top of visual representations using conventional policy networks composed of CNNs and MLPs in the WidowX setup and diffusion policy [72] in the Spot setup. During testing, we vary the robot position for Door Opening and Drawer Opening, randomize the object position for Pick-and-Place, and randomize both the object positions and object types in Toy-Microwave Cooking. Full experimental settings, including details about the number of collected demonstrations, the interface used to collect them, and the policy architecture, are available in the Appendix (Sec. D.3).
@@ 5 What Makes Visual Representations Good for Robot Learning?
Darcet et al. [24] analyzed the norm of spatial tokens in vision Transformers and found that high-norm outlier tokens are detrimental to vision task performance. Following this, we investigate whether a similar phenomenon arises in visual representations for robot learning. We inspect the feature norms of Theia with different teacher combinations and baseline models evaluated in Section 4.2, and their corresponding performance on the MuJoCo subset tasks. We sample 1% of the MuJoCo task training set and calculate the L2-norm of each spatial token after encoding images. We measure the entropy of the feature norm distribution across all samples and patches per model and use it as a quantitative metric. To calculate the entropy of the distribution, we first discretize the distribution by a histogram. Then, we use the following formula to obtain the entropy $H=-\sum_{i}(p_{i}*\log(p_{i}))$, where $p_{i}$ is the probability of each discretized bin.
We confirm that similar outlier tokens also appear in VC-1 corresponding to the image patches that are not task-relevant, shown in the visualizations of feature norms on the right of Figure 7. In contrast, Theia has very few or no outlier tokens, and the tokens with higher norms are more task-relevant even though Theia-representations are not trained on these robot images. In our quantitative analysis (Figure 7, left), we divide the models into distilled and regular based on the observation that distilled models generally have higher entropy (fewer outliers, Figure 4(c) in Darcet et al. [24]). We find that there is a strong correlation (R=0.943) between entropy and robot learning performance among regular models, and a high correlation (R=0.638) among distilled models. We hypothesize that spatial token representations with high entropy (better feature diversity) encode more information that aids policy learning, while less diverse representations (low entropy) may hinder it. In the Appendix G, we discuss the results of other quantitative measurements, including feature similarity and PCA-explained variance ratios, where no strong correlations are found.
@@ Appendix A Theia Model Architecture
Feature Normalization.
Formally, the normalization is:
$\tilde{h(\mathbf{x}_{j})}_{c}=\frac{h(\mathbf{x}_{j})_{c}-\mu_{c}}{\sigma_{c}},~{}\mu_{c}=\frac{1}{N}\sum_{j}{h(\mathbf{x}_{j})_{c}},~{}\sigma_{c}=\sqrt{\frac{\sum_{j}{(h(\mathbf{x}_{j})_{c}-\mu)^{2}}}{N}},$ (2)
where $c$ is the channel index for the feature, $j$ is the index of image sample, and $N$ is the number of samples in ImageNet.
@@ D.2 CortexBench
For ImageNav task, we use the provided policy network from VC-1 [20] without modification. The policy network is composed by a compression-layer (a CNN layer) to convert the spatial feature map into a vector representation, followed by a 2-layer LSTM. Details are available in Appendix A.2 in [20].
@@ D.3 Real World Robot Learning
Table 9: Number of demonstrations and data collection method for the real-world experiments.
Task Door-Opening Pick-and-Place Toy-Microwave Cooking Drawer Opening
# Demos 48 63 101 50
Method Teleoperation Teleoperation Teleoperation Scripted Policy
@@ D.3.1 WidowX Arm Experiments
We train a behavior-cloning policy for each of the four evaluated setups and for each evaluated baseline (see Table 2 and Figure 3); the training hyperparameters are shown in Table 10. To train each of the three tasks performed with the WidowX robot, i.e., Door Opening, Pick-and-Place, and Toy-Microwave Cooking, we collected human demonstrations by teleoperating the robot with a Virtual Reality (VR) controller using the setup introduced in  [59]; the number of collected demonstrations for each of the tasks is reported in Table 9.
@@
In this task the robot has to open a fridge door in a toy-kitchen setup; we identify two stages to evaluate the task’s success: Open and Fully Open (see Figure 3). We place the robot in front of the fridge and collect 63 demonstrations to train the behavioral cloning policy. We vary the height (z-axis) of the robot base between 40-46cm, and the position (x-axis, parallel to the toy kitchen) of the robot base; samples from the initial state distribution of the demonstrations are shown in Figure 8. At inference time, we vary the height of the robot base among $\{40,42,43,44,46\}$ cm, and select between 5 randomly-picked positions along the x-axis (for all policies). For each robot position, we evaluate the policy twice, for a total of 50 runs.
Figure 8: Samples from the initial state distribution of the collected demonstrations for fridge door-opening (top), pick-and-place (middle), and toy-microwave cooking (bottom). For each of these tasks, the factors of variation are, respectively: the robot height and position with respect to the door, the initial pose of the cup, and the initial position of the objects and object types.
@@
In this task, the robot has to pick up a pink cup from a toy-sink and drop it into a drying rack located on the left of the sink. We collected 48 demonstrations to train the policy, where we varied the initial pose of the objects, as shown in Figure 8. During evaluation, we vary the cup’s starting position amongst a total of 10 positions, of which 8 positions are equally distributed about the perimeter, and 2 are in the center of the sink. We also roughly vary the direction of the cup handle towards the left or the right. In total, we evaluate this task for 20 runs. There are two key stages for which we measure the success rate: picking up the cup and successfully releasing it into the drying rack.
@@
In this task, the robot has to pick up an object from within the pot on the stovetop, putting the object into a toy-microwave, and closing the microwave. In each test, we initialize the environment with the microwave door open. In this task, we collected 100 demonstrations across 10 different toy-food objects (10 demonstrations per object) with randomized object positions; examples from the initial state distributions of the collected demonstrations can be seen in Figure 8. During evaluation, we test 40 runs on 10 seen, in-distribution objects (4 runs per object), and 10 runs on 5 unseen, out-of-distribution objects (2 each), for a total of 50 runs. Furthermore, we vary the position of the pot that holds the object. We point out that this task is characterized by a longer horizon with respect to the other two; in fact, it has three different steps: 1. picking up the object, 2. placing the object into the microwave, 3. and closing the microwave. In addition, the policy needs to generalize to the different poses of the objects on the stove-top and to out-of-distribution objects. The longer horizon and the variety of objects in this task make it particularly challenging, so using frozen visual encoders was not effective (0% success rate). However, with fine-tuning, the policies performed reasonably well.
Figure 9: Samples from the initial state distribution of the collected demonstrations for the Spot drawer opening experiments. For this task, the factors of variation are the position and orientation of the robot with respect to the drawer.
@@ D.3.2 Spot Experiments
In this task, the robot has to open the top drawer of a cabinet. The policy receives color images from a single forward facing camera mounted on the body of the Spot. A fiducial marker is used to enable the robot to walk to a random position and orientation for each trial. Samples from the distribution of random initial states are shown in Figure 9. The starting locations vary by ±$\pm 5$ cm in x and y and ±$\pm 0.2$ radians in orientation. 50 successful demonstrations were collected using a scripted policy and we evaluate each policy for 20 trials. A trial is considered successful if the drawer is opened at least 10 cm. After each trial, a scripted policy uses the fiducial marker to reset the environment and the robot moves to a new random location.